Please help me figure out this problem (x-2)x-(x-2)
Answer by stanbon(75887)

(x-2)x-(x-2)
=x^2-2x-x+2
=x^2-3x+2
OR
(x-2)x-(x-2)
=(x-2)(x-1)
=x^2-3x+2
Cheers,
stan H.